OpenStudy (anonymous):

Jake had grades of 92,90,85 and 88 on five math test . What grade must he get on the 5th test to have an average of 90 ?

OpenStudy (tamtoan):

if x is the fifth test score, the average = total score divide 5 tests, try that

OpenStudy (anonymous):

Do you know how to fine out the average?

OpenStudy (anonymous):

Average = Sum of observations / total no. of observations Take the unknown grade as x and then work it out yourself! :)

OpenStudy (anonymous):

(92+90+85+88+x)/5=90 (355+x)/5=90 355+x=450 x = 95 He must score a 95 on the next test to get an average of 90 in the class